OpenStudy (dreamyemo):

square root of 11(square root of 2 - square root of 7)

OpenStudy (anonymous):

\[\sqrt{11}\left( \sqrt{2}-\sqrt{7} \right)=\sqrt{22}-\sqrt{77}\]